OpenStudy (anonymous):

(3x+15)+(2x-10) = 90

OpenStudy (kc_kennylau):

1. Remove the parentheses 2. Combine like terms together 3. Throw the things unrelated to x to the RHS 4. Divide both sides by 5.

OpenStudy (anonymous):

x=17 if that's what your looking for

OpenStudy (haseeb96):

3x+15 +2x-10 =90 5x+5=90 5x=90-5 5x=85 x=85/5 x=17 this is the correct answer